When 15.0 mL of 2.50 M iron (II) sulfate are allowed to react with 15.0 mL of 1.00 M
sodium sulfide, a precipitate is formed. What is the maximum mass of precipitate that
can be formed from this reaction?
1
Expert's answer
2020-06-11T10:28:05-0400

FeSO4 + Na2S = Na2SO4 + FeS.

2.5 x 15/1000 = 0.0375.

15 x 1/1000 = 0.015.

(56+32) x 0.015 = 1.32 g.